OpenStudy (anonymous):

Parabola Help Need to put into general form, y = a(x - h)2 + k I have 0,0 and 10,0 for the x- intercept, and 5,3 for the vertex, the parabola goes upward. How do I put it into general form? Thanks

OpenStudy (anonymous):

vertex is (h, k), so h = 5 and k = 3. and then plug in 0,0 and 10,0 for x and y and solve the system of 2 equations

OpenStudy (anonymous):

so y=a(10-5)2+3 Is that right? or am I not done?

OpenStudy (anonymous):

no you have to do 0 = a(10-5)^2 + 3, so 25a = -3, so a = -3/25. Then the general form is y = -3/25(x-5)^2 + 3

OpenStudy (anonymous):

okay thanks